# FwNet-ECA: Facilitating Window Attention with Global Receptive Fields through Fourier Filtering Operations

### v1

@@ Abstract
Windowed attention mechanisms were introduced to mitigate the issue of excessive computation inherent in global attention mechanisms. However, In this paper, we present FwNet-ECA, a novel method that utilizes Fourier transforms paired with learnable weight matrices to enhance the spectral features of images. This strategy facilitates inter-window connectivity, thereby maximizing the receptive field. Additionally, we incorporate the Efficient Channel Attention (ECA) module to improve communication between different channels. Instead of relying on physically shifted windows, our approach leverages frequency domain enhancement to implicitly bridge information across spatial regions. We validate our model on the iCartoonFace dataset and conduct downstream tasks on ImageNet, demonstrating that our model achieves lower parameter counts and computational overheads compared to shifted window approaches, while maintaining competitive accuracy. This work offers a more efficient and effective alternative for leveraging attention mechanisms in visual processing tasks, alleviating the challenges associated with windowed attention models. Code is available at https://github.com/qingxiaoli/FwNet-ECA
@@ 1 Introduction
However, the pursuit of capturing global spatial relationships by Self Attention models incurs a quadratic increase in computational complexity with respect to image resolution. A common approach to mitigate this involves encoding neighboring pixels into a single token. Despite this, the model still demands considerable parameters and computation, with limitation particularly evident in tasks such as image segmentation and object detection. Liu et al. [1]addressed this by partitioning images into smaller pixel regions and employing windowed attention to reduce computational load. To overcome the restricted receptive field of local attention, they further proposed shifted window attention for cross-window information exchange. Nevertheless, this method solely enhances communication among the chosen windows.
In this work, we use filter enhancement operations to establish a global receptive field for window attention, thereby avoiding the use of shifted window attention. The purpose of using filtering operations is to enhance or suppress certain frequencies, allowing the filter-enhanced window attention to more clearly identify where to focus its attention. Following each Filtering Enhancement, we append adjacent channel attention to bolster inter-channel interactions, termed FwNet-ECA. Unlike shifted window attention, the Filtering Enhancement layer directly facilitates connections among tokens across the entire feature map. Leveraging the Fast Fourier Transform (FFT) in two dimensions, the computational complexity of the Filter Enhancement layer is reduced to O(NlogN).
@@ 2 Related Works
The advent of Transformers in natural language processing has catalyzed a paradigm shift in computer vision, leading to the exploration of ViT[11]. Pioneered by Dosovitskiy et al. [2], ViT treat images as sequences of patches, enabling more flexible modeling of global dependencies without the need for convolutional operations. This approach achieved impressive results in image classification but was inherently constrained by high computational requirements and large model sizes[12]. Subsequent improvements aimed to tackle these issues through architectural refinements, often reintroducing some level of inductive bias. A notable example is the Swin Transformer[1], which curtails computational complexity via localized attention mechanisms. To facilitate interaction between local regions, it introduces shifted window attention, albeit at the cost of increased computational inefficiency and parameter count due to elaborate masking operations, which are also limited in fostering global communication among windows. This work seeks a simpler structure for inter-window connectivity.
In the visual domain, the Fourier transform, particularly the Discrete Fourier Transform (DFT), appears in many convolutional neural networks. With CNNs demonstrating substantial advancements in visual recognition and analysis, various strategies have emerged to optimize and innovate for specific visual tasks. One such method involves exploiting frequency domain representations through DFT, thereby enhancing task performance by leveraging frequency features[13][14]. Another concurrent approach leverages the convolution theorem in conjunction with FFT to expedite CNN computations[15]. Some other works are based on CNN to construct the interdependence between channels in the frequency domain[16]. Within Transformer architectures, networks like FNet[17] and GFNet[18] have replaced Self Attention mechanisms with filtering operations for both CV and NLP tasks.
Regarding channel attention, SENet[19] initiated the concept of channel-wise attention, enabling networks to dynamically adjust feature channel weights through its SE(Squeeze-and-Excitation) module, significantly boosting model performance. Subsequent innovations, such as ECA-Net[20] and CBAM[21], not only improved computational efficiency but also fortified feature processing capabilities by integrating spatial attention, finding broad application across diverse deep learning tasks[22][23]. In this work, we added an ECA(Efficient Channel Attention) module to the main structure of the model to construct inter channel attention.
@@
  • Different purposes: Although similar to GFNet in method, the purpose of building FwNet-ECA is to find a simple and efficient method for information interaction between windows and solve the limitations of windows.
@@ 4 Experiment
Across all model training processes, we employed a series of data augmentation techniques to improve the generalization ability of our models. Specifically, these included resizing images to a fixed size ($(img\_size,img\_size)$), applying a RandomResizedCrop with scale ranging from $0.15$ to $1.0$ and aspect ratio between $3/4$ and $4/3$, using RandomHorizontalFlip for horizontal flipping, and utilizing ColorJitter to randomly adjust the brightness, contrast, and saturation within a range of $0.4$. Additionally, images were converted into tensors and normalized using mean values of $[0.485,0.456,0.406]$ and standard deviation values of $[0.229,0.224,0.225]$. Our models were initialized with pre-trained weights from ImageNet1k. We compared the computational cost, parameter counts, and Top-1 accuracy of different models.
@@ 4.3 Analysis
In the following analysis, it should be noted that for aesthetic reasons, the curves were smoothed when drawing Fig.4 5 6. To analyze the changes in the number of parameters and computational complexity of filter enhancement, self-attention, and windowed self-attention with respect to the feature map size and token dimensions. We conducted a detailed analysis of the Filter Enhancement layer within a hierarchical model structure, using the same patch size and initial token dimension for embedding operations. In a hierarchical structure, the token dimension is doubled after each stage, and the height and width of the feature map are halved, leading to a rapid increase in dimensionality. The change in parameter count with increasing token dimensions is illustrated in Fig. 4. Since the windowed self-attention does not change the number of parameters compared to self-attention, we have only plotted the curve for the variation of the number of parameters for self-attention. Initially, the parameter count for Filter Enhancement exceeds that of Self Attention; however, as the token dimension increases, the parameter count for Self Attention grows exponentially, whereas the parameter count for Filter Enhancement decreases. This suggests that Filter Enhancement is more sensitive to image resolution or token quantity, while Self Attention is more sensitive to token dimension.
@@ 5 Conclusion
The proposed FwNet-ECA achieves information exchange between different windows in window attention through Filter Enhancement operations based on Fourier Transform, reducing the computational cost of inter-window communication from $O(N^{2})$ to $O(Nlog(N))$. This approach enables the model to achieve the largest receptive field with reduced parameters and computational overhead. Additionally, it establishes communication between channels through ECA modules. Experimental results confirm the effectiveness of the Filter Enhancement layer and ECA modules. Furthermore, our model demonstrates certain advantages in fine-grained classification tasks.

### v2

# FwNet-ECA: A Classification Model Enhancing Window Attention with Global Receptive Fields via Fourier Filtering Operations

@@ Abstract
Windowed attention mechanisms were introduced to mitigate the issue of excessive computation inherent in global attention mechanisms. In this paper, we present FwNet-ECA, a novel method that utilizes Fourier transforms paired with learnable weight matrices to enhance the spectral features of images. This method establishes a global receptive field through Filter Enhancement and avoids the use of moving window attention. Additionally, we incorporate the Efficient Channel Attention (ECA) module to improve communication between different channels. Instead of relying on physically shifted windows, our approach leverages frequency domain enhancement to implicitly bridge information across spatial regions. We validate our model on the iCartoonFace dataset and conduct downstream tasks on ImageNet, demonstrating that our model achieves lower parameter counts and computational overheads compared to shifted window approaches, while maintaining competitive accuracy. Furthermore, our visualization operations clearly demonstrated that the Filter Enhancement technique achieves greater effectiveness in the model’s shallow layers, where feature maps are relatively larger. This work offers a more efficient and effective alternative for leveraging attention mechanisms in visual processing tasks, alleviating the challenges associated with windowed attention models. Code is available at https://github.com/qingxiaoli/FwNet-ECA
@@ 1 Introduction
However, the pursuit of capturing global spatial relationships by Self Attention models incurs a quadratic increase in computational complexity with respect to image resolution. A common approach to mitigate this involves encoding neighboring pixels into a single token. Despite this, the model still demands considerable parameters and computation, with limitation particularly evident in tasks such as image segmentation and object detection. Liu et al. [1]addressed this by partitioning images into smaller pixel regions and employing windowed attention to reduce computational load. To overcome the restricted receptive field of local attention, they further proposed shifted window attention for cross-window information exchange. However, this method only enhances the connections between partial windows and lacks a global receptive field. Moreover, the implementation of shifted windows involves complex masking operations and tensor shape transformations, making it relatively complicated. We aim to achieve a similar effect in a simpler way.
In this work, we use filter enhancement operations to establish a global receptive field for window attention, thereby avoiding the use of shifted window attention. The purpose of using filtering operations is to enhance or suppress certain frequencies, allowing the window attention to more clearly identify where to focus its attention. Following each Filtering Enhancement, we append adjacent channel attention to bolster inter-channel interactions, termed FwNet-ECA. Unlike the moving window attention, the filter enhancement layer directly obtains a global receptive field, and the computation is simple, requiring only Fourier transforms and element-wise matrix multiplication. This is easy to implement based on the PyTorch library. Leveraging the Fast Fourier Transform (FFT) in two dimensions, the computational complexity of the Filter Enhancement layer is reduced to O(NlogN).
@@
We further substantiate the effectiveness of the filter enhancement operation, especially in the shallow layers of models where feature maps are relatively large, through a series of visualization experiments. This finding demonstrates that employing filter enhancement techniques in the initial stages of a model can significantly improve the quality of feature extraction, thereby enhancing overall model performance. This effect is particularly beneficial for tasks that require capturing more macroscopic and global information from input images.
@@ 2 Related Works
The advent of Transformers in natural language processing has catalyzed a paradigm shift in computer vision, leading to the exploration of ViT[11]. Pioneered by Dosovitskiy et al. [2], ViT treat images as sequences of patches, enabling more flexible modeling of global dependencies without the need for convolutional operations. This approach achieved impressive results in image classification but was inherently constrained by high computational requirements and large model sizes[12]. Subsequent improvements aimed to tackle these issues through architectural refinements, often reintroducing some level of inductive bias. A notable example is the Swin Transformer[1], which curtails computational complexity via localized attention mechanisms. To facilitate interaction between different windows, it introduces shifted window attention, which is a very clever method but is complex in operation. In this work, we aim to propose a simpler and more intuitive method.
In the visual domain, the Fourier transform, particularly the Discrete Fourier Transform (DFT), appears in many convolutional neural networks. With CNNs demonstrating substantial advancements in visual recognition and analysis, various strategies have emerged to optimize and innovate for specific visual tasks. One such method involves exploiting frequency domain representations through DFT, thereby enhancing task performance by leveraging frequency features[13][14]. Another concurrent approach leverages the convolution theorem in conjunction with FFT to expedite CNN computations[15]. Within Transformer architectures, networks like FNet[16] and GFNet[17] have replaced Self Attention mechanisms with filtering operations for both CV and NLP tasks.
Regarding channel attention, SENet[18] initiated the concept of channel-wise attention, enabling networks to dynamically adjust feature channel weights through its SE(Squeeze-and-Excitation) module, significantly boosting model performance. Subsequent innovations, such as ECA-Net[19] and CBAM[20], not only improved computational efficiency but also fortified feature processing capabilities by integrating spatial attention. In this work, we added an ECA(Efficient Channel Attention) module to the main structure of the model to construct inter channel attention.
@@
  • Different purposes: Although the method is similar to GFNet, the purpose of constructing FwNet-ECA is to find a simple and efficient method to replace the more complex moving window attention.
@@ 4 Experiment
Across all model training processes, we employed a series of data augmentation techniques to improve the generalization ability of our models. Specifically, these included resizing images to a fixed size $(224,224)$, applying a RandomResizedCrop with scale ranging from $0.15$ to $1.0$ and aspect ratio between $3/4$ and $4/3$, using RandomHorizontalFlip for horizontal flipping, and utilizing ColorJitter to randomly adjust the brightness, contrast, and saturation within a range of $0.4$. Additionally, images were converted into tensors and normalized using mean values of $[0.485,0.456,0.406]$ and standard deviation values of $[0.229,0.224,0.225]$. Our models were initialized with pre-trained weights from ImageNet1k. We compared the computational cost, parameter counts, and Top-1 accuracy of different models.
@@ 4.2 Imagenet downstream tasks
4.3 Visualization
To more clearly demonstrate how our filter enhancement boosts the establishment of a global receptive field to assist window attention, we visualized the responses of images from the filter enhancement modules and shifted window attention modules at different stages in both FwNet-ECA and Swin Transformer. For this visualization, we employed three types of Class Activation Mapping (CAM) methods: GradCAM[28], XGradCAM[29], LayerCAM[30], and the outputs of each method were averaged to obtain the Integrated CAM. The rationale behind choosing these methods is that while GradCAM is suitable for rapid implementation and broadly applicable, it may underperform in complex scenarios. XGradCAM, on the other hand, improves accuracy and interpretability through normalization, making it ideal for tasks requiring high-precision activation maps. LayerCAM, which integrates feature maps from multiple layers, is particularly apt for complex scenes and multi-scale analysis. Concerning model selection, we utilized a tiny-sized model with pre-training conducted on the iCartoonFace dataset. As shown in 5, stages 1 and 4 each contain only one filter enhancement or shifted window attention block, whereas in stage 3, we selected the middle module for examination.
Figure 4: Visual Comparison of Filter Enhancement Operations vs. Shifted Window Attention in the Initial Stage of the Model
We chose this image for visualization because it possesses high resolution and distinct features, such as unique lines and an exaggerated artistic style, making it ideal for observation through visualization techniques. As shown in GradCAM of Fig. 4, our frequency enhancement operation in stage 1 demonstrates a grasp on the overall outline of the character and key features like the curve of the character’s abdomen. XGradCAM highlights the main focus of the model by effectively reducing noise; however, our model does not show significant responses in images generated by LayerCAM. This is likely due to the relatively uniform gradients in the shallow layers of our model, whereas LayerCAM calculates gradient information for each pixel, aiming to capture finer-grained local responses. This also explains why the visualization of shifted window attention shows no response in GradCAM and XGradCAM but reveals a small area of interest in LayerCAM. Although the gradients of the shifted window attention are minimal, there is a notable change in a localized region at the lower left part of the character’s abdomen in the image.
Figure 5: Visual Comparison of Filter Enhancement Operations vs. Shifted Window Attention in the Middle Stage of the Model
Figure 6: Visual Comparison of Filter Enhancement Operations vs. Shifted Window Attention in the Last Stage of the Model
As we’ve conducted visualizations in the intermediate and final stages of the model, as shown in Fig. 5 6 respectively, it’s evident that the filter enhancement responses generally cover a larger area and respond at more positions across the image compared to the shifted window approach. The response locations of the shifted window method show a clear blocky distribution, which is especially noticeable in the final stage. In these stages, it can be seen that the shifted window responses are concentrated in the central region and a square area in the upper left of the image.
This observation suggests that while filter enhancement tends to activate over wider areas and potentially capture broader or more diverse features within an image, the shifted window attention mechanism focuses its responses more narrowly, possibly honing in on specific features or areas of interest with higher precision. This distinction highlights the different strategies these methods employ for feature extraction and their implications on the model’s understanding of the input images.
@@ 4.4 Analysis
In the following analysis, it should be noted that for aesthetic reasons, the curves were smoothed when drawing Fig.7 8 9. To analyze the changes in the number of parameters and computational complexity of filter enhancement, self-attention, and windowed self-attention with respect to the feature map size and token dimensions. We conducted a detailed analysis of the Filter Enhancement layer within a hierarchical model structure, using the same patch size and initial token dimension for embedding operations. In a hierarchical structure, the token dimension is doubled after each stage, and the height and width of the feature map are halved, leading to a rapid increase in dimensionality. The change in parameter count with increasing token dimensions is illustrated in Fig. 7. Since the windowed self-attention does not change the number of parameters compared to self-attention, we have only plotted the curve for the variation of the number of parameters for self-attention. Initially, the parameter count for Filter Enhancement exceeds that of Self Attention; however, as the token dimension increases, the parameter count for Self Attention grows exponentially, whereas the parameter count for Filter Enhancement decreases. This suggests that Filter Enhancement is more sensitive to image resolution or token quantity, while Self Attention is more sensitive to token dimension.
@@
4.5 Deficiencies and prospects
The limitation of our model lies in the fact that for hierarchical models, when the feature map resolution is small, the filter enhancement method may not yield satisfactory results. This is because the frequency resolution of the feature map in the frequency domain is also low at such times, making it difficult to precisely distinguish between different frequency components, which affects the filtering effect. Meanwhile, low-resolution feature maps inherently lack high-frequency details, and information may be further lost after filtering, leading to subpar performance.
However, precisely because of this, we are hopeful about the performance of this method on high-resolution images, as they are likely to avoid several of the aforementioned issues. However, due to computational power limitations, we have not been able to conduct experiments at higher resolutions.
@@ 5 Conclusion
The proposed FwNet-ECA achieves information exchange between different windows in window attention through Filter Enhancement operations based on Fourier Transform, reducing the computational cost of inter-window communication from $O(N^{2})$ to $O(Nlog(N))$. This approach enables the model to achieve the largest receptive field with reduced parameters and computational overhead. Additionally, it establishes communication between channels through ECA modules. Experimental results confirm the effectiveness of the Filter Enhancement layer and ECA modules, and through visualization methods, it is proved that the filtering enhancement can indeed obtain a better receptive field than the shifted window attention. Furthermore, our model demonstrates certain advantages in fine-grained classification tasks.